Ian reported that the Preservation Dept. would start to get part time employees in a week. Both voucher and work study will work for the department. Ian reported that most of the moldy items from the Alexander Library project have been returned. A note for remediation will be added to the Bib record. Approximately 500 items need barcodes and brief records.
For the SMLR Union Conference Procedures Project, Paul reported that 70 titles have been completed, with 80 titles remaining. The Doc Ref State Project is currently on hold, with staff waiting for word from Elizabeth L. as to when the project can continue. 40 monographs are remaining. For the Doc Ref Foreign Project, 20 titles are remaining. The Dana Inventory Project is currently on hold, due to problems with the equipment. There has been no inventory performed at Dana since the 2nd week of August. There is a small amount of BibCat work left to work on. At LSM, still working with Neera and Sal on BibCat errors and errors found in the collections. Paul will continue with the monthly meeting with Neera and Sal. Paul created a PowerPoint for title changes, which he has shared with Jessica at Alexander and Neera at LSM. Jessica at Alexander working on pre-cat material and searching the monographic series titles from the 16 boxes of materials that Lourdes brought down. 8 boxes have been processed so far. Paul met with Elizabeth L. to continue discussing serials to monographs problems.
Sal discussed his visit to Dana Library, where he found out that the binding of current periodicals has been stopped due to possible funding problems. With Karen leaving soon, Sal will test to see if it is possible to do receiving at LSM with just two receivers. Alexander staff could lend a receiver if there is a backlog. Sal continues to work on the backlog of monographs. Sal reported that Bill is training part time employees, with 15-20 currently working at LSM.
Yuhwei continues to work on the duplicate print list. He is also waiting for publishers to get back to him with information. Yuhwei reported that Linda has been training part time employees; where at Alexander this semester they currently have 20-25 work study employees. For EAL, Alexander DTS staff are receiving current issues for them. He has scheduled a meeting with Ying, Kayo, and Connie on 9/28 to discuss the backlog of gifts at EAL. Yuhwei has scheduled 10 hours a week for part time employees to work at EAL. Yuhwei discussed serial control records for online only titles. These control records will not be for checkin purposes but for notification purposes through the claims feature. Yuhwei distributed a document concerning online only title serial controls and the EAL workflow for the group to review.
Andy reported on the work with the Douglass Folio Collection, where the items have been internally relocated online and are in the process of being labeled. For the Math Library Reference Collection, the items have been transferred online to the Annex and will be shipped soon. The training session for phase two of the Serials Inventory Project for the NBL Branches has been scheduled.
In the past BIBCAT staff was creating brief records for titles received through packages. Rebecca reported that the Receiving Coordinators would now be doing that. Rebecca discussed the amount of cancellations that Rob has been involved in. The transfer to Harrossowitz has meant a lot of coordination between staff in the department. Rebecca asked the group to be aware of E-only periodical titles and related serial control records.
Jamie reported that all items with accompanying materials at TAS would be sent to LSM DTS for final processing. This change in the process will address any possible processing errors that may come up with this material.
Gracemary discussed the CDC meeting that was held that morning. The Serials vendor has been changed to Harrassowitz. There has been some difficulty in communicating with the US service rep, so DTS staff have been contacting the German service rep. DTS staff will be trained by a Harrassowitz service rep on 9/28. With the change, order control and serial control records need to be change. Claims are more difficult, but they will need to be changed before January 1st. With the change, DTS will try to implement EDI (Electronic Data Interchange) with help from Systems for invoicing.
Gracemary discussed the LIS meeting that she attended. The two new Preservation users were discussed at the meeting. She also discussed the status of the request to replace “TECHSR” for DTS use with another use. The name of the user, “LOCALMAINT” was decided at the group, which should be coming soon. Gracemary also discussed a problem that was discovered with the Federal Register in the DOC collection. It was found that the item records were changed from shelved by title to SUDOC at one time. This could be due to a Marcive upload problem, where overlays of existing records could change item lines for non-duplicates. Depending on the solution to this problem, this may lead to DTS staff examining these records. Finally, Gracemary announced that Bob Hosh is in charge of the Government Processing Group.
Melody reported that the LSM weeding project is running smoothly. Over 32,000 items have been pulled from the LSM collections, with over 35,000 items overall processed by LSM DTS. Melody will work with Neera to update the procedures for library wide weeding, based on experiences from the LSM weeding project.
A document discussing the relevant workflows for materials that crossover between DTS and Preservation was distributed to the group for review. Access Services staff will continue to send material in need of repair directly to the DTS hubs, after checking the item out to the appropriate CS User. Staff at DTS will evaluate the items and decide whether DTS staff will continue with the work or if the item needs to be sent to Preservation. If the item will be sent to Preservation, DTS staff will discharge the item from the CS user and charge to the appropriate Preservation user. Item sent to Preservation will include moldy items, brittle material, pamphlet binding, customized pockets/housing, and material in need of reback/recase only. All DTS processing should be completed before the item is sent to Preservation.
The group discussed the needs of the two DTS hubs when it comes to work study employees. It is very important that team leaders communicate their needs for part time employees, so that the department can maximize the efficiency of the part time employees that they do have. The group discussed the new check distribution procedures.
The group discussed the upcoming upgrade to 3.1 and the upcoming change to the J client for Workflows. Gracemary will send out the forms for testing for the upcoming upgrade to 3.1. The upgrade to 3.1 will happen in December; so all testing must be completed this semester. The group needs to look at the new J client together and consider what the change to this new software will entail. All staff will move to the J client at once, to make sure that everyone is looking at the same information.
Next meeting: October 19, 2006
Next recorder: Gracemary Smulewitz
Back up recorder: Melody Tomaszewicz